# Preserving Multilingual Quality While Tuning Query Encoder on English Only

### v1

@@ 4 Conclusion
We considered tuning the query part of a dual encoder starting from a high quality multilingual embedding model, and using English-only samples in the tuning. We found that multilingual qualities are quite stable in many scenarios of the tuning, and can be not only preserved but improved. We explain this by speculating that most of the transformer, except the embedding block, depends weakly on multiple languages. We also think of this as a particular case of a more general pattern: tuning a certain model quality, if done carefully enough, can improve the related qualities too. In our setting this allows the simple and resource-light adjustment of multilingual embeddings for specific query types or domains.
5 Limitations
Our consideration here is limited to starting with a single high quality multilingual embedding model, and tuning it (on English-only samples) as a query encoder. While this setup is good for understanding and convenient for adjusting an existing model, it would be natural to follow this up by considering a pre-trained multilingual dual encoder which is already asymmetric from the start.
For our illustration we used the state of the art multilingual model intfloat/multilingual-e5-small, and also, for comparison, repeated the same observations for the sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 model. To gain a better understanding of the observed behaviours, it would be good to investigate more multilingual models.
@@ 5 Limitations
We used MSMARCO triplets for tuning; for evaluation we used a set aside part of MSMARCO triplets, and ARXIV in two variations, and XNLI. The motivation was that the MSMARCO evaluation part must show improvement (after tuning), ARXIV must verify the robustness of the improvement on a very different kind of texts, and XNLI must reveal the effect of the English-only driven improvement on multilingual qualities. That said, the evaluations can be extended to more datasets.
Appendix A Triplets from ARXIV subset
We make use of triplets created from arxiv because this provides our evaluation with a very different kind of text (compared to MSMARCO), and thus allows us to judge robustness of the improvement. The procedure of making the triplets is simple, and it may be more appropriate to provide a recipe rather than duplicate already existing data: arxiv-metadata-oai-snapshot.jsonl from mirror of arxiv (version 173). As outlined in Section 2.2, the triplets are created in the following steps:
@@ Appendix A Triplets from ARXIV subset
    Identify all arxiv categories with a maximum size of 10K articles (i.e. arxiv preprints).
@@
    Select in the same order as given all samples (articles) that have at least one of the categories identified above. This is the subset of arxiv to deal with: manageably small, yet diverse. Keep the order of the samples in the subset, and iterate over them in this order as following:
@@
    For each sample $A$ in the subset, pick a “random” sample $B$ (random seed 0, python version 3.8). If the picked up $B$ coincides with $A$, select instead the sample next after $B$. If no next, select the first sample. Create a triplet either as (1) the title of sample $A$ as the anchor, its abstract as positive, and abstract of sample $B$ as negative, or (2) the first sentence of the abstract of sample $A$ as the anchor, the rest of the abstract as positive, and the abstract of sample $B$ (first sentence excluded) as negative.
The difference made by the “randomness” of the selection of negatives is negligible, but if exact reproducibility on arxiv is wanted, fixing the random seed to 0 in python 3.8 should do it.
learning msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
rate cdcdcdc+/- d+/- c+/- d+/-
1e-8 0.73 0.94 -0.35 0.00 0.50 -0.44 0/0 0/0 0/0 0/0
2e-8 6.50 7.62 4.71 4.75 2.76 5.96 222/0 218/0 208/1 171/5
3e-8 6.65 7.57 4.56 4.57 3.51 5.96 222/0 217/0 202/1 170/6
4e-8 6.66 7.58 4.31 4.62 2.76 6.18 222/0 218/0 207/1 174/5
5e-8 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
6e-8 7.14 8.74 4.63 5.10 3.26 7.28 222/0 218/0 204/2 168/13
7e-8 9.11 11.35 5.58 5.84 3.01 8.83 223/0 216/3 193/17 140/52
8e-8 8.75 10.98 5.75 6.50 4.51 9.05 222/0 217/3 186/20 131/61
9e-8 9.01 11.26 5.78 6.17 4.01 8.83 222/0 216/3 182/25 129/67
1e-7 9.26 11.38 5.38 6.12 3.51 9.49 222/0 214/3 188/18 132/63
Table 3: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, batch size 14, margin 0.1 and different learning rates (first column). Values that did not pass the two-tailed test are shown in gray.
size cdcdcdc+/- d+/- c+/- d+/-
7 6.85 7.57 3.86 4.44 3.01 6.62 221/0 217/0 198/2 163/13
14 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
28 7.45 9.47 4.46 5.13 4.01 8.39 222/0 219/1 196/6 145/27
56 6.51 7.33 3.46 4.14 2.51 5.74 221/0 217/0 202/1 168/6
112 4.63 4.78 1.84 2.16 2.51 2.21 212/0 203/0 196/0 155/1
size cdcdcdc+/- d+/- c+/- d+/-
7 6.76 7.99 3.69 4.72 3.26 6.18 221/0 216/0 177/7 119/31
14 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
28 8.50 10.18 5.03 6.12 2.76 8.17 222/0 218/3 197/9 147/36
56 7.42 9.12 3.86 4.47 3.76 8.17 223/0 220/0 200/2 160/21
112 9.50 11.84 -1.54 -4.42 -15.54 -14.13 175/35 146/65 91/117 32/175
margin msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
cdcdcdc+/- d+/- c+/- d+/-
0.05 7.40 9.32 3.71 3.83 2.76 5.30 225/0 223/0 201/4 169/21
0.10 7.38 8.95 4.91 4.90 3.26 7.73 222/0 218/0 201/2 163/17
0.15 6.90 7.78 6.00 5.79 4.26 7.51 213/0 204/3 176/7 125/33
0.20 6.40 7.23 7.32 6.80 5.01 8.39 175/6 161/14 132/32 94/67
0.25 5.09 5.70 7.42 7.11 5.51 7.73 127/27 116/42 108/55 76/84
0.30 3.92 4.23 6.50 6.35 4.51 6.40 108/40 99/52 97/62 79/85
0.35 3.76 3.80 4.53 4.54 4.01 3.53 80/40 74/49 72/57 65/75
0.40 2.77 3.10 5.06 5.41 4.51 5.30 69/72 67/69 64/80 61/94
0.45 2.59 2.78 4.68 4.85 4.51 4.64 54/81 59/77 58/83 56/98
0.50 2.94 2.41 2.99 2.44 2.51 1.77 30/23 31/27 53/19 48/26
Table 6: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14 and different triplet loss margins (first column). Values that did not pass the two-tailed test are shown in gray.
to stop cdcdcdc+/- d+/- c+/- d+/-
5 6.50 7.60 4.09 4.60 2.26 6.62 222/0 217/0 208/1 174/5
10 7.38 8.95 4.91 4.90 3.26 7.73 222/0 218/0 201/2 163/17
15 8.93 10.98 5.83 6.07 2.51 7.73 222/0 217/3 191/17 140/51
@@
Table 8: Evaluations of the $L12$ query model tuned on MSMARCO as described in Section 2.3. The rows are in the order of increased freezing (at tuning): from no freezing (top row) to freezing everything up to the last transformer block (the last row); the last block is $B11$. The $emb.base$ model has only the first three layers of the embedding block frozen (tokens, positions, token-types). The $emb$ model has the full embedding block frozen. The other notations: $B0$ is the full first transformer block; $B0$-$5$ are the first 6 blocks; the extensions $a$, $i$, $od$ (for $B0$) correspond to the layers $attention$, $intermediate$ and $output.dense$ of the block. The columns $c\%$ and $d\%$ show the improvement (in percents) relative to the original $L12$ model by evaluation with cosine ($c$) and distance ($d$). The columns $c$+/- and $d$+/- show how many language pairs are improved ($+$) or worsened (--) by statistically significant change of the number of errors, measured with cosine ($c$) or distance ($d$). Values that did not pass two-tailed test are gray-colored.
@@
learning msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
rate cdcdcdc+/- d+/- c+/- d+/-
1e-8 1.85 2.15 0.76 -1.17 3.75 0.63 55/15 0/15 131/15 0/16
2e-8 4.81 5.18 1.77 -3.98 8.32 1.22 199/15 41/25 196/15 6/59
3e-8 4.70 5.11 1.94 -4.03 8.74 1.29 198/15 39/25 195/15 6/64
4e-8 7.22 8.03 2.23 -8.98 12.21 -1.13 203/15 50/41 207/15 20/134
5e-8 7.28 8.04 2.28 -9.23 12.4 -1.03 200/15 47/47 206/15 20/143
6e-8 7.16 7.91 2.55 -9.07 12.69 -0.98 203/15 52/41 207/15 20/135
7e-8 9.41 10.72 0.88 -20.77 10.94 -12.1 174/28 20/174 201/16 16/185
8e-8 9.45 10.69 1.23 -20.97 11.64 -12.03 178/28 25/164 201/15 16/182
9e-8 9.29 10.57 2.03 -19.69 12.88 -9.92 187/25 28/144 203/15 17/178
1e-7 9.40 10.74 1.84 -20.44 12.43 -10.69 180/27 28/160 202/15 17/177
@@
Table 9: Evaluations of the $L12$ query encoder tuned with a frozen embedding block, batch size 14, margin 0.1 and different learning rates (first column). Values that did not pass the two-tailed test are shown in gray.
@@
margin msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
cdcdcdc+/- d+/- c+/- d+/-
0.05 9.00 10.16 1.09 -22.22 10.23 -13.81 180/28 27/159 205/15 21/171
0.10 7.28 8.04 2.28 -9.23 12.4 -1.03 200/15 47/47 206/15 20/143
0.15 9.90 11.34 1.11 -19.64 11.98 -10.18 164/29 19/177 193/20 15/195
0.20 9.71 11.16 1.58 -16.94 12.63 -6.99 163/29 20/171 193/22 15/198
0.25 9.30 10.75 1.39 -13.92 12.83 -4.2 163/27 16/160 183/22 14/205
0.30 9.62 11.1 1.21 -13.82 12.6 -3.94 125/33 13/180 157/31 14/209

### v3

@@ 4 Conclusion
We considered tuning the query part of a dual encoder starting from a high quality multilingual embedding model, and using English-only samples in the tuning. We found that multilingual qualities are quite stable in many scenarios of the tuning, and can be not only preserved but improved. We explain this by speculating that most of the transformer, except the embedding block, depends weakly on multiple languages. We think of this as a particular case of a general pattern: tuning a certain model quality, if done carefully enough (adiabatic tuning), can also retain or even improve the related (but not targeted by tuning) qualities. This allows a resource-light adjustment of multilingual embeddings for a specific query type or domain, even a narrow domain (Appendix K).
Limitations
Our considerations here are limited to starting with a single high quality multilingual embedding model, and tuning it (on English-only samples) as a query encoder. While this setup is good for our understanding and convenient for adjusting an existing model, it would be natural to follow this up by considering a pre-trained multilingual dual encoder which is already asymmetric from the start.
For our illustration we used the state of the art multilingual model intfloat/multilingual-e5-small, and also, for comparison, repeated the same observations for the sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 model. We also repeated some of observations on monolingual model intfloat/e5-small-v2 - the tuning improved its rudimentary multilingual properties as well. Still, to gain a better understanding of the observed behaviors, it would be interesting to investigate more multilingual models.
@@
We used MSMARCO triplets for tuning; we also verified some observations for models tuned on ARXIV-based subsets limited to a category (math, physics or cs, Appendix K). For evaluation we used a set aside part of MSMARCO triplets, and ARXIV in two variations, and XNLI. The motivation was that the MSMARCO evaluation part must show improvement (after tuning), ARXIV must verify the robustness of the improvement on a very different kind of texts (jargon-heavy), and XNLI must reveal the effect of the English-only driven improvement on multilingual qualities. We also confirmed the tuning gains on SQUAD and HotpotQA (both of which are quite different from MSMARCO). That said, the evaluations can be extended to even more datasets.
More research could be helpful in understanding and identifying the range of adiabatic tuning.
Appendix A Usage of MSMARCO Triplets
The MSMARCO dataset consists of 499184 samples, with each sample being a tuple given as (query, positives, negatives). The “positives” are the correct answers to the query, and the “negatives” are semantically similar, but incorrect answers. For most samples there is only one positive, but many negatives. For our tuning we simply select the very first positive and the very first negative. Thus, each sample gives one triplet (anchor, positive, negative) for contrastive learning, where the query is taken as an anchor.
We keep the first 487983 samples (or 34856 batches if each batch is 14 triplets) for tuning, leaving the next 4200 samples (300 batches) for validation, and the last 7000 samples for evaluation. During evaluation we create all possible triplets from the 7000 samples, using all positives and negatives; this makes 357642 evaluation triplets.
Almost half of MSMARCO samples have the maximal number of negatives (65), and for evaluation shown in Table 2 we use a more difficult version ’MSMARCO 65 negatives’, with all samples with less than 65 negatives filtered out.
Appendix B ARXIV Dataset for Triplets
B.1 Dataset arxiv-negatives
We made two flavors of evaluation arxiv triplets from this arxiv subset. In the first version, the anchor is the title, the positive is the corresponding abstract, and the negative is another random abstract. In the second version the anchor is the first sentence of the ’positive’ abstract, the positive is the rest of the abstract, and the negative is a similar piece (first sentence excluded) of the ’negative’ abstract.
We make use of triplets created from arxiv because this provides our evaluation with a very different kind of text (compared to MSMARCO), and thus allows us to judge robustness of the improvement. For convenience and reproducibility of creating triplets of different levels of difficulty, we made a dataset arxiv-negatives [^10] https://huggingface.co/datasets/primer-ai/arxiv-negatives. The dataset consists of 253140 samples, each sample is a tuple of two elements:
@@ B.1 Dataset arxiv-negatives
    (1) An ARXIV paper metadata, including its Id, title and abstract and categories.
@@
    (2) List of 21 Ids of other ARXIV papers. The first 20 Ids are the papers that are ’closest’ to the above paper, and sorted from the most to the least similar; the last 21st Id is an Id of a randomly selected paper (not coinciding with Id of the above paper).
Thus, we have 21 versions of picking up negatives for triplets, from the most difficult to the easiest (the last one, of the random selection).
For example, to create triplets of difficulty 14, for each paper given by the first tuple element, we pick up a paper corresponding to 14th Id given in the second tuple element. From the first paper we can create query and positive, and from the second paper - negative. Through this work we used two flavours:
  1. 1.
    ’Title’: Query is the title of the first paper, and positive is its abstract; negative is the abstract of the second paper.
  2. 2.
    ’First’: Query is the first sentence of the abstract of the first paper; positive is the rest of the abstract; negative is the abstract of the second paper, with its first sentence deleted.
B.2 How is it created?
The above dataset is created from the mirror of arxiv [^11] https://huggingface.co/datasets/arxiv-community/arxiv_dataset [^12] https://www.kaggle.com/datasets/Cornell-University/arxiv (version 173) arxiv-metadata-oai-snapshot.jsonl through the following steps:
  1. 1.
    Identified all arxiv categories with a maximum size of 10K papers (i.e. arxiv preprints).
  2. 2.
    Selected all papers that have at least one of the categories identified above. This is the subset of arxiv to deal with: manageably small, yet diverse.
@@ B.2 How is it created?
    For each paper: (1) Sort its categories by size, from smaller to larger. (2) Find all other papers that have the closest match by the categories (the closest match is the longest consecutive list of matched categories, starting from the first one). (3) Of the found papers, select 20 closest by Jensen-Shannon distance between the paragraphs, and sort them by the distance. If there were less than 20 papers, fill to 20 by the last one. (4) Add randomly selected paper as 21st.
Of the total 253140 samples, in 213156 samples (84.2%) all the first 20 negatives are different (which means that not less than 20 papers happen to have the same closest match by categories).
Appendix C SQUAD
For using SQUAD dataset, we identified (for each query) the given paragraph sentences containing an answer to the query as positives, and the rest of the sentences as negatives. We left samples having at least 1 positive and 1 negative. On average there is 1.3 positives and 4.2 negatives per a query. For evaluation shown in Table 2 we combined train, validation and test subsets. The results are given also for a version ’SQUAD min 5’, in which we have filtered out queries that had less than 5 candidate sentences.
Appendix D HotpotQA
For using HotpotQA, we combined its train and dev subsets. For each query (’question’) both train and dev subsets contain on average 9.95 passages, of which 2 are always positives. For evaluation shown in Table 2 we filtered out queries that had less than 10 passages, and split the dataset into ’easy’, ’medium’ and ’hard’ subsets accordingly to the HotpotQA labels of the difficulty of the samples.
Appendix E Tuning
Unless specified otherwise, we tune a dual encoder by contrastive learning in the following simple regime:
  1. 1.
    The text encoder is fully frozen; the frozen parts of the query encoder are specified.
  2. 2.
    The batch size is 14, the learning rate is 5e-8 and the contrastive learning margin is $0.1$. The loss is defined by the triple margin loss.
  3. 3.
    There are $1000$ batches per epoch, i.e. 14000 samples per epoch.
  4. 4.
    Stopping occurs after $10$ consecutive non-improvement epochs. The improvement is measured on the validation subset after each epoch. The model is considered to be improved if (on the validation subset) both the loss and the count of errors have decreased.
  5. 5.
    The AdamW optimizer is used.
Changing this default regime is considered in Appendixes LM.
Appendix F XNLI
The XNLI dataset consists of pairs of sentences which are human-labeled as entailment, neutral or contradiction. The test subset (which we use) contains 1670 pairs for each of these labels and each sentence is presented in 15 languages: [’ar’, ’bg’, ’de’, ’el’, ’en’, ’es’, ’fr’, ’hi’, ’ru’, ’sw’, ’th’, ’tr’, ’ur’, ’vi’, ’zh’]. We use 225 versions of the pairs, because each sentence of the pair can be in any of the 15 languages. At evaluation the first sentence serves as the query (the embedding is taken by the query model), and the second one as the text. We expect that the sentences of an entailment pair should be closer to each other than the sentences of any neutral pair, or of any contradiction pair. Whenever this does not happen, we count this as an error.
Appendix G Performance of Untuned Query Encoder
To establish a baseline before any fine-tuning, and to ensure our evaluation is not too easy, we measure the errors of the original $E5$ model on the data described in Section 2.3 and show the results in Table 3. We also measure the errors of $L12$ and of $E5e$ a more recent monolingual (English) model.
data Evaluation $E5$ $L12$ $E5e$
MM N tot 357642
PND (cos) 4.7% 15.1% 4.6%
PND (dist) 4.8% 15.4% 4.5%
ARX-F N tot 253140
PND (cos) 1.6% 4.9% 3.1%
PND (dist) 1.6% 6.7% 3.5%
ARX-T N tot 253140
PND (cos) 0.2% 1.4% 0.2%
PND (dist) 0.2% 1.7% 0.2%
XNLI N total 2788900
PND e-n (cos) 10.8% 10.2% 15.9%
PND e-c (cos) 10.0% 7.2% 15.3%
PND e-n (dist) 10.5% 10.1% 15.9%
PND e-c (dist) 9.6% 7.8% 15.4%
Table 3: The count of errors for the original untuned models $E5$, $L12$ and $E5e$, on the datasets noted in the first column: MM - MSMARCO test 7000 samples (357642 triplets, see Section 2.2 and Appendix A); ARX-F - arxiv-first, the arxiv subset with the abstract’s first sentence as an anchor; ARX-T - arxiv-title, the arxiv subset with the title as an anchor; XNLI - XNLI test subset providing 1670x1670=2788900 comparisons of entailment pairs vs neutral pairs (and the same amount of entailment pairs vs contradiction pairs). For XNLI the errors are averaged over 225 (15x15) language-language versions, and shown as percent of $Ntotal$. The evaluation is done using cosine similarity or euclidean distance similarity (cos or dist in second column).
The count of errors on the triplets (MSMARCO, ARXIV) is straightforward: it is an error when a positive is not closer than a negative to the anchor of the triplet. On XNLI we sum up the error count over all language-language pairs and divide the sum by the number (255=15x15) of such pairs. This averaged error is shown as a percentage of the total (2788900) comparisons; each comparison here is either a comparison of an entailment-labeled sample with a neutral-labeled sample (entail-neutral in the table) or a comparison of an entailment-labeled sample with a contradiction-labeled sample (entail-contr in the table). An error was counted whenever the sentences of an entailment sample happened to be farther from each other than the sentences of a neutral (or contradiction) sample. Separately for each pair of languages PND is shown in Figures 34 for cosine similarity measure. The distance measure gives results visually almost undistinguishable.
Figure 3: PND of embedding models on XNLI entailment-neutral comparisons assessed by cosine.
Figure 4: PND of embedding models on XNLI entailment-contradiction comparisons assessed by cosine.
The amount of errors in Table 3 and in Figures 34 is reasonable enough to consider how tuning would affect them. The smallest counts are the counts of positive-negative discrepancies of $E5$ and $E5e$ on ARX-T (apparently, a title makes an easier ’query’ than the first sentence of an abstract). These counts are 309 and 420 for the cosine similarity (the row ARX-T PND (cos)), and 453 and 580 for the distance similarity (the row ARX-T PND (dist)).
Notice that $L12$ has far worse PND on English data (MSMARCO and ARXIV). The English-only model $E5e$, as expected, performs worse than multilingual models $E5$ and $L12$ on multilingual XNLI, but its PND is still far below $50\%$, because there is much similarity between some of the languages.
Appendix H Gains on ARXIV for Different Levels of Difficulty
Figure 5: Errors and improvements on arxiv-negatives dataset of different level of difficulty. The “easiest” dataset is a random selection of negatives from the same data used through this work in evaluations. In (a), we show the fraction of errors done by the original E5 model (for comparison, see Table 3). In (b), we show the improvement after tuning the query encoder on MSMARCO, with ‘default’ settings, i.e. learning rate 5e-8, batch size 14, margin 0.1 and frozen embedding block. Values that did not pass the two-tailed test (Appendix I) are shown with open markers.
Through the paper we used the easiest version of triplets in the dataset arxiv-negatives, the version that uses randomly selected negatives. Here in Figure 5 we show, for comparison, the fraction of the errors that untuned original $E5$ embeddings do on the other levels of difficulty, and also the corresponding improvements (by Equation 1) after tuning the query encoder on the MSMARCO with frozen embedding block and our default settings (Section 2.3). The statistical significance of the improvements in Figure 5 is estimated as explained in Appendix I.
The difficulty of intentionally close negatives is much harder, but Figure 5 still shows that performance on ARXIV was mostly improved. We used the easiest triplets version for our evaluations through the paper because it more distinctly indicated the trends in the improvements.
Appendix I Significance Test
In Table 1, Figure 2 and through the paper we use two-proportion $Z$-test, pooled for $H_{0}:p_{1}=p_{2}$. We are comparing the number of errors original $n_{0}$ and improved $n_{1}$, having the total $N$ (the totals can be seen in Table 3); a total is the same for original and improved version. We deem the difference to be significant if $|Z|>Z_{c}$ where
$Z=\frac{p_{1}-p_{0}}{\sqrt{\frac{1}{2}P(1-P)N}}$ (2)
with $p_{0}=n_{0}/N$, $p_{1}=n_{1}/N$ and $P=\frac{1}{2}(n_{0}+n_{1})/N$. We used $Z_{c}=1.96$, which is a critical value corresponding to probability 0.975.
Notice that in our examples the values $N$ are typically very large. And the improvements we report, according to Equation 1, are relative, not absolute values.
Appendix J Encoder L12 with Frozen Layers
The changes in cross-lingual qualities corresponding to the third row (emb, frozen embedding block) of Table 4 are shown in comparison with $E5$ and $E5e$ embeddings in Figures 6 and 7. Note that $E5e$ is not a multilingual embedding model. Having a worse start as a multilingual embedding model, $E5e$ also gets much weaker improvements of its multilingual qualities; it is consistent with our understanding of adiabatic tunings (Section 3.2).
@@ Appendix I Significance Test
Table 4: Evaluations of the $L12$ query model tuned on MSMARCO as described in Section 2.3. The notations are as in Table 1.
Figure 6: Improvement of $E5$, $L12$ and $E5e$ on XNLI entailment-neutral comparisons assessed by cosine.
Figure 7: Improvement of $E5$, $L12$ and $E5e$ on XNLI entailment-contradiction comparisons assessed by cosine.
Appendix K Narrow-Domain Query Encoder
So far we observed that tuning the query encoder on data of a certain style (MSMARCO dataset) could preserve (or even improve) the encoder qualities which are not targeted by the tuning task, especially if we tune with a frozen embedding layer and low learning rate. Here we provide observations using more specialized datasets, based on arxiv-first (arxiv-first is described in Section 2.2 and Appendix B):
  1. 1.
    ARXIV-math: uses only documents with at least one category which has the prefix "math."
  2. 2.
    ARXIV-physics: As above, but with "physics." as the prefix
  3. 3.
    ARXIV-cs: As above, but with "cs." as the prefix
$E5$ tuned on these narrow datasets using our ‘default’ regime (Section 2.3) with frozen embedding block mostly improves the PND (positive-negatives discrepancy fraction) as shown in Table 5. The improvements of these narrow-tuned encoders on individual language pairs, assessed by cosine, are shown in Figures 8 and 9.
msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
model cdcdcdc+/- d+/- c+/- d+/-
E5-math 0.04 0.45 54.71 52.85 50.38 53.64 218/0 209/0 177/0 133/0
E5-physics 0.16 0.57 19.05 18.64 21.8 20.97 162/0 102/0 31/0 2/0
E5-cs 0.18 0.55 23.32 23.63 25.56 26.49 205/0 136/0 51/0 8/8
Table 5: Evaluations of the $E5$ query encoder tuned on ARXIV-math, ARXIV-physics or ARXIV-cs with a frozen embedding block, batch size 14, margin 0.1 and learning rate 5e-8. When evaluated on ARXIV (columns arxiv-first and arxiv-title) the samples with category of the model (the first column) are excluded from the evaluation data.
Figure 8: Improvement of narrow-tuned encoders on XNLI entailment-neutral comparisons assessed by cosine.
Figure 9: Improvement of narrow-tuned encoders on XNLI entailment-contradiction comparisons assessed by cosine.
Appendix L Learning Rate
In Figure 2 we have shown how the improvements of the $E5$ model depend on the learning rate. Here in Figure 10 we compare similar data for $L12$ and $E5e$ as well as a particular instance of $E5$ when both the query and text encoder are subject to tuning (as two independent encoders, with the same starting point) with the embedding block frozen in both encoders. The data confirm that while higher learning rates are not yet overtuning and still give higher gains on the test subset (of MSMARCO), it is the lower learning rates that better preserve and even improve those pretrained qualities which are not the goal of tuning.
Figure 10: Improvement of various models and tuning configurations on the English-only datasets (MSMARCO and ARXIV) in the left column and XNLI in the right column. Values that did not pass the two-tailed test (Appendix I) are shown with open markers. (a) Evaluations of the $E5$-full dual encoder after both encoders were tuned with a frozen embedding block, batch size 14 and margin 0.1. (b) Evaluations of the $L12$ query encoder tuned with a frozen embedding block, batch size 14 and margin 0.1. (c) Evaluations of the $E5e$ query encoder tuned with a frozen embedding block, batch size 14 and margin 0.1.
Appendix M Tuning Regime
M.1 Learning rate and batch size
M.1.1 Scaling rule
@@
batch learning msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
size rate cdcdcdc+/- d+/- c+/- d+/-
7 2.5e-8 6.62 7.67 4.48 5.06 3.26 6.18 221/0 216/0 201/2 160/13
14 5.0e-8 7.30 8.82 4.90 5.39 3.51 7.73 222/0 217/1 201/2 158/20
28 1.0e-7 8.36 10.31 5.71 6.75 3.26 7.95 222/0 218/3 177/20 128/58
56 2.0e-7 8.32 10.54 5.39 6.75 2.76 7.51 222/0 217/3 193/14 141/42
112 4.0e-7 8.46 10.36 5.24 6.00 3.26 8.39 221/0 216/3 197/8 147/35
@@
Table 6: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, margin 0.1 and 14000 samples per epoch. Linear scaling rule of learning rate with batch size. Values that did not pass the two-tailed test are shown in gray.
batch learning msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
size rate cdcdcdc+/- d+/- c+/- d+/-
7 3.54e-8 6.80 7.84 4.43 4.76 2.26 6.18 221/0 216/0 192/2 148/17
14 5.00e-8 7.30 8.82 4.90 5.39 3.51 7.73 222/0 217/1 201/2 158/20
28 7.07e-8 7.16 9.10 4.70 5.57 4.01 7.95 222/0 217/1 192/4 140/32
56 1.00e-7 7.32 8.56 4.70 5.16 3.01 6.40 221/0 217/0 204/2 169/13
112 1.41e-7 7.25 8.55 5.24 4.91 4.01 7.51 222/0 217/0 202/2 166/16
Table 7: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, margin 0.1 and 14000 samples per epoch. Square root scaling rule of learning rate with batch size. Values that did not pass the two-tailed test are shown in gray.
The learning rate is usually set with consideration to the batch size; it can be proportional to the batch size (linear scaling rule), or proportional to square root of the batch size (square root scaling rule) Krizhevsky (2014); Goyal et al. (2018); Hoffer et al. (2018). We show the evaluation results for these scaling rules in Tables 6 and  7. While there is no essential wins in scaling batch size and learning rate up or down, the square root rule seems more reasonable in keeping the evaluation results approximately the same while increasing the batch size.
Regardless of the overall behavior of scaling the batch size and learning rate together, we have to verify that our default batch size 14 is a good fit for our default learning rate $5e$-$8$. For this reason, a simple change of batch size, without altering learning rate, is considered in Appendix M.1.2; the tables 8 and 9 show that our ‘default’ batch size is reasonable. The corresponding data for $L12$ are in Appendix M.1.3.
M.1.2 Encoder E5 and the batch size
size cdcdcdc+/- d+/- c+/- d+/-
7 6.85 7.57 4.55 4.78 3.01 6.62 221/0 217/0 198/2 163/13
14 7.30 8.82 4.90 5.39 3.51 7.73 222/0 217/1 201/2 158/20
28 7.45 9.47 5.14 5.46 4.01 8.39 222/0 219/1 196/6 145/27
56 6.51 7.33 4.16 4.48 2.51 5.74 221/0 217/0 202/1 168/6
112 4.63 4.78 2.55 2.50 2.51 2.21 212/0 203/0 196/0 155/1
size cdcdcdc+/- d+/- c+/- d+/-
7 6.76 7.99 4.38 5.06 3.26 6.18 221/0 216/0 177/7 119/31
14 7.30 8.82 4.90 5.39 3.51 7.73 222/0 217/1 201/2 158/20
28 8.50 10.18 5.71 6.45 2.76 8.17 222/0 218/3 197/9 147/36
56 7.42 9.12 4.55 4.81 3.76 8.17 223/0 220/0 200/2 160/21
112 9.50 11.84 -0.82 -4.05 -15.54 -14.13 175/35 146/65 91/117 32/175
In Table 8 we show results for batch sizes 7, 14, 28, 56 and 112, while keeping the number of samples per epoch the same (14000). The row with batch 14 here coincides with the values for learning rate 5e-8 in Figure 2, and with the row for the frozen embedding block in Table 1. The results for all batch sizes are similar. Tuning with the higher batch size of 112 is a bit ‘safer’ for languages, not degrading any language pair when evaluated by cosine measure, and degrading only one language pair (for entailment vs. contradiction) when evaluated by distance measure. This comes at the price of lower gains on MSMARCO and ARXIV.
@@
Table 9 shows what happens if the number of batches per epoch (1000) is kept the same, rather than the number of samples. In this setting the larger batch size of 112 leads to a less frequent validation (by MSMARCO validation subset) at tuning and, effectively, to later and less reasonable stopping. This results in higher gains on MSMARCO test subset, but in far worse results on ARXIV and XNLI.
M.1.3 Encoder L12 and the batch size
Dependency of $L12$ tuning on the batch size is shown in Table 10 (number of samples per epoch is 14000) and in Table 11 (number of batches per epoch is 1000). Observations are somewhat similar to $E5$ (Appendix M.1.2), except that generally $L12$ performs not as good as $E5$ and the batch size 7 turns out bad for $L12$.
M.2 Weight decay
@@
weight msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
decay cdcdcdc+/- d+/- c+/- d+/-
100 2.77 1.88 -2.47 -1.06 4.01 1.32 84/104 80/99 90/96 71/92
50 5.39 5.00 0.49 2.12 3.26 2.43 120/51 121/41 140/51 133/46
10 7.08 8.36 3.54 5.11 4.26 6.62 222/0 217/0 201/3 160/18
5 7.88 9.84 4.97 5.87 3.01 7.95 222/0 216/2 189/15 144/36
1 7.26 8.70 4.72 5.11 3.01 7.06 222/0 218/0 202/2 164/16
0.5 7.28 8.78 4.87 5.24 3.01 7.06 221/0 218/0 202/2 163/18
0.1 7.32 8.56 4.70 5.16 3.01 6.40 221/0 217/0 204/2 169/13
0.05 7.32 8.56 4.70 5.16 3.01 6.40 221/0 217/0 204/2 169/13